OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3.

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (anonymous):

i don't have the option of y-y1 =m(x-x1)

OpenStudy (anonymous):

y+2=-3(x-7) or y=19-3x

OpenStudy (anonymous):

can you help with another one kristin848?

OpenStudy (anonymous):

sure, I'll see what I can do :)

OpenStudy (anonymous):

Choose the correct procedure used to graph the equation 3x + 9y = -18 using the intercepts method. Solve the equation for y to get a y-intercept of (0, -2) and a slope of -1/3 . Plot the y-intercept, count the slope to get another point and draw the line that passes through the two points. Choose two different values; plug them into the equation to find the corresponding y-values. Plot the two points and draw the line that passes through the two points. Let x equal zero and solve for y to find a y-intercept of (0, -2). Let y equal zero and solve for x to find an x-intercept of (-6, 0). Plot the two points and draw a line passing through the two points. Solve the equation for y to find an intercept of (-2, 0); solve the equation for x to find a y-intercept of (0, -6). Plot the two points and draw a line passing through the two points.

OpenStudy (anonymous):